OpenStudy (anonymous):

evaluate limit as x->0 (sinx/x)^(cot(x)^2)

geerky42 (geerky42):

This?\[\Large \lim_{x \rightarrow 0} \left( \dfrac{\sin (x)}{x} \right)^{\cot^2(x)}\]

OpenStudy (anonymous):

lim sin(x) x->0 -------- = 1 x is the theorem. Now, cot^2x = cos^2x / sin^2x=1/0=undefined. So I think the limit doesn't exist.

geerky42 (geerky42):

Limit does exist

OpenStudy (anonymous):

Yes, geerky42, that's the one. I know the answer, but I need to actually show the process of getting it and wolfram alpha is saying no step-by-step solution is available.

geerky42 (geerky42):

No idea how to approach it... anything you learned recently that would help us with this problem?

OpenStudy (anonymous):

we are working on indeterminate limits such as 0/0, inf/inf, 1^inf, etc. I feel like I have the first few steps right, but it seems there are some trig identities being used and I just can't figure out which one to use when

OpenStudy (anonymous):

have you learned taylor series yet?

OpenStudy (anonymous):

sin(x)~=x-x^3/6 (taylor series) (sin(x)/x)^(cot^2(x)) = (1-x^2/6)^(1/sin(x)^2-1) = (1-x^2/6)^(1/(x-x^3/6)^2-1) = (1-x^2/6)^(1/x^2-1) = (1-x^2/6)^(1/x^2) let u=-6/x^2. as x->0, u->infinity = ((1+1/u)^u)^(-1/6) = e^(-1/6)

OpenStudy (anonymous):

we have not done taylor series yet

OpenStudy (anonymous):

Did u do it?

OpenStudy (anonymous):

no, still haven't figured it out yet

OpenStudy (anonymous):

Wait here,for some minutes..

OpenStudy (anonymous):

I have tried it about 7 different ways, but no luck yet

OpenStudy (anonymous):

Is the answer is 1.

OpenStudy (anonymous):

no, e^(-1/6)

OpenStudy (anonymous):

what about lhopital? :3

OpenStudy (anonymous):

yes, I've used it, but there is much more to it with the exponentital stuff

OpenStudy (anonymous):

I haven't. Mind if I give it a shot? Does it give a dead end? ^.^

OpenStudy (anonymous):

\[\Large L=\lim_{x \rightarrow 0} \left( \dfrac{\sin (x)}{x} \right)^{\cot^2(x)}\] \[\Large \ln L=\lim_{x \rightarrow 0} \ \ln\left( \dfrac{\sin (x)}{x} \right)^{\cot^2(x)}\] \[\Large \ln L=\lim_{x \rightarrow 0} \ [\cot^2(x)]\ln\left( \dfrac{\sin (x)}{x} \right)\]

OpenStudy (anonymous):

And then we have (for nothing more than my personal preference :D) \[\Large \ln L=\lim_{x \rightarrow 0} \left[\frac{\ln\left(\frac{\sin(x)}{x}\right)}{\tan^2(x)}\right]\] So far, so good, right, @ajnelson23 ? ^.^

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

can't seem to make the right side equal -1/6

OpenStudy (anonymous):

your answers... they make me feel like i'm flying towards a bloody wall.... or at least a wall that will be bloody after I crash into it XD

OpenStudy (anonymous):

yep...been working on this for like 2 days looking all over at rules I've never heard of to see if I can make some sense of this

OpenStudy (anonymous):

anyway, let's use lhopital on the right side, differentiate both top and bottom of the fraction, we get... \[\Large \ln L=\lim_{x \rightarrow 0} \left[\frac{x\csc(x)\cdot\frac{x\cos(x)-\sin(x)}{x^2}}{2\sec^2(x)\tan(x)}\right]\]

OpenStudy (anonymous):

actually\[\Large \ln L=\lim_{x \rightarrow 0} \left[\frac{\csc(x)\cdot\frac{x\cos(x)-\sin(x)}{x }}{2\sec^2(x)\tan(x)}\right]\]

OpenStudy (anonymous):

yep, got that far

OpenStudy (anonymous):

And it looks terrible XD

OpenStudy (anonymous):

Let's see... \[\Large \ln L=\lim_{x \rightarrow 0} \left[\frac{\csc(x)\left[x\cos(x) - \sin(x)\right]}{2x\sec^2(x)\tan(x)}\right]\]

OpenStudy (experimentx):

Don't use L'hopital after it ... multiply by x^2 at the top and ... and bottom of x cos(x) - sin(x) / x^3 ... evaluat ethe limits individually.

OpenStudy (anonymous):

wait, what?

OpenStudy (anonymous):

Still playing with stuff. I'm a bit lost @experimentX :D \[\Large \ln L=\lim_{x \rightarrow 0} \left[\frac{x\cos(x) - \sin(x)}{2x\sec(x)\tan^2(x)}\right]\]

OpenStudy (experimentx):

Just evaluate this limit .. http://www.wolframalpha.com/input/?i=Limit+x-%3E+0++%28x+cos%28x%29+-+sin%28x%29%29+%2F+x%5E3 rest is easy.

OpenStudy (experimentx):

use L'hopital here 3 times ... or you can do it without using L'hopital or power series.

OpenStudy (experimentx):

but it get's complicated here :(

OpenStudy (anonymous):

I do hate it when I have to L'Hopital more than once :(

OpenStudy (anonymous):

so multiply the whole thing by (x^2/x^2) and do l'hospital's again?

OpenStudy (experimentx):

yes ... that's what I meant.

OpenStudy (experimentx):

@PeterPan you like doing it without using L'hopital ?? there is a way ... just change x -> 3x but this will get complicated.

OpenStudy (anonymous):

how does multiplying (x^2/x^2) times peter pan's limit up there equal the limit you put in wolfram alpha? I am not seeing how to get from one to the other?

OpenStudy (experimentx):

hmm what you mean?? the other is just of the form ... just check it x^2/2 sin(x)^2

OpenStudy (experimentx):

that cos thing will go to 1 as x->0

OpenStudy (anonymous):

ok, to be more specific, which section above should I multiply times (x^2/x^2)?

OpenStudy (experimentx):

\[ \Large \ln L=\lim_{x \rightarrow 0} \left[\frac{x^2\csc(x)}{2\sec^2(x)\tan(x)}\right] \frac{x\cos(x)-\sin(x)}{x^3 } \] .To get this

OpenStudy (experimentx):

Evaluate the left limit and the right limit separately ... on left side, you get 1/2 on right side you get -1/3 hence answer is e^(-1/6) best way to do this is to use Taylor series as done above.

OpenStudy (anonymous):

ok, let me try to work it

OpenStudy (experimentx):

the left side is easy ... the right side is hard.

OpenStudy (anonymous):

I got!!! Thanks so much!

OpenStudy (experimentx):

very good.